# Phys4D: Fine-Grained Physics-Consistent 4D Modeling from Video Diffusion

### 03 Mar 2026 v1

@@ 1 Introduction
Recent video generation models (OpenAI, 2024; Google DeepMind, 2024; Polyak et al., 2024) and current generative world models(Chen et al., 2025c; Yu et al., 2024b, 2023; Huang et al., 2025b) have achieved impressive visual realism and large-scale physics. Yet their success does not imply physical understanding: most are trained to match appearances rather than to model underlying physical state and dynamics. As a result, they frequently fail at fine-grained physics, producing locally inconsistent geometry, unstable object motion, and non-causal dynamics over time (Singer et al., 2022; Zhou et al., 2024). A faithful world model should instead represent how the environment evolves, requiring physics-consistent 4D representations.
A key challenge in physics-consistent 4D world modeling is the lack of scalable supervision for fine-grained physical dynamics. Learning such models requires dense, temporally aligned signals of geometry and motion, which are extremely difficult to obtain from real-world data at scale. Existing 4D datasets(Wen et al., 2025a; Zhai et al., 2024) therefore tend to be appearance-rich but physics-poor, often dominated by camera motion. To address this limitation, we leverage the diversity and controllability of physics-based simulation to obtain large amounts of dynamic video data with accurate ground-truth geometry and motion, providing supervision that is otherwise inaccessible.
@@
To evaluate whether a model behaves as a physics-consistent 4D world model, we introduce a comprehensive evaluation protocol that combines content-level physics assessment and 4D world-level physics analysis. Specifically, in addition to established video-based physics benchmarks such as PhysicsIQ(Motamed et al., 2025b), which evaluate physical plausibility from visual and semantic cues, we propose a 4D evaluation suite that probes geometry and motion world representations through trajectory-based analysis. Together, these complementary evaluations assess physical plausibility from what a model shows to how the underlying world evolves, providing a holistic diagnosis of physics-consistent 4D modeling beyond appearance and reconstruction quality.
@@ 2 Related Work
Recent video generation models(OpenAI, 2024; Google DeepMind, 2024) achieve strong visual realism and temporal coherence, but are largely optimized for appearance rather than physical world modeling. Several works attempt to introduce physical cues into video generation, such as modeling object dynamics(Zhang et al., 2024; Aira et al., 2024) or action-conditioned motion(Chang et al., 2024; Yuan et al., 2025), and a number of benchmarks(Motamed et al., 2025b; Meng et al., 2025; Bansal et al., 2024b) evaluate physical plausibility based on visual or semantic signals. While these approaches can improve or assess appearance-level physical realism, they operate primarily in image space and do not evaluate whether a model maintains coherent geometry, motion, or a consistent physical state over time. In contrast, our work focuses on physics consistency at the level of a 4D world, motivating geometry- and time-aware evaluation beyond appearance-centric video physics assessment. Detailed discussion please refer to Appendix G.1 G.2 G.3
@@
A growing body of work explores incorporating physical priors into generative models, including post hoc physical simulation after generation (Stomakhin et al., 2013b; Chen et al., 2025a; Xie et al., 2024), simulation-guided generation with predefined dynamics (Yuan et al., 2023; Liu and others, 2024a; Li et al., 2025), and learning physical priors from data (Li et al., 2024; Xu et al., 2024; Yang et al., 2025a). These methods typically focus on object-level or short-horizon dynamics, rely on specialized 3D representations, or remain optimization-based, limiting scalability and generalization. In contrast, our work leverages pretrained video diffusion models and treats fine-grained physics as a property of an evolving 4D world. For a detailed discussion, please refer to Appendix G.4
@@ 4 Simulation Data Collection
We present a physics-based simulation pipeline designed to address the limited physical fidelity and diversity of existing training data for 4D world modeling. In contrast to prior datasets that focus on narrow physical regimes or specific simulation paradigms (Huang et al., 2025a; Chang et al., 2024), Phys4D supports large-scale data generation with diverse object types, rich physical interactions, and accurate geometry and motion supervision. Figure 2 provides an overview of the simulation pipeline and the range of supported object categories.
@@ 6.1 Video-Based Physics Evaluation
We evaluate on Physics-IQ (Motamed et al., 2025a), a benchmark designed to assess physical understanding in generative video models. The benchmark comprises 198 test scenarios covering diverse physical phenomena. Each scenario includes three camera viewpoints and two natural variations, yielding 1,188 test videos.
@@
We apply Phys4D to three backbone: WAN2.2-5b (Wan et al., 2025), CogVideoX-5b (Yang et al., 2025b), and Open-Sora V1.2 (Zheng et al., 2024). For evaluation, we generate 5-second videos conditioned on the switch frame following the Physics-IQ protocol.
@@ F.1 Flow Matching for Video Generation
which induces a Gaussian conditional distribution $\mathbf{z}_{t}\sim\mathcal{N}(t\mathbf{z}_{1},(1-t)^{2}\mathbf{I})$. Notably, flow matching is not restricted to Gaussian endpoints or linear paths and can be generalized to interpolate between arbitrary distributions (Albergo and Vanden-Eijnden, 2023; Liu et al., 2023). This flexibility has recently enabled efficient large-scale generative modeling, including applications to high-dimensional and structured data such as videos, where flow-based formulations offer favorable trade-offs between sampling efficiency and temporal coherence (Xu et al., 2022; Lin and others, 2024b; Peng and others, 2025).
@@ F.4 Policy Optimization with PPO
where $\hat{A}_{t}$ denotes the estimated advantage. This objective stabilizes policy updates and has been shown effective for fine-tuning diffusion- and flow-based generative models (Black and others, 2023; Ren and others, 2024; Liu and others, 2025).
@@ G.1 Large-Scale Video Generation and World Models
Recent large-scale video generation models, such as Sora (OpenAI, 2024), Veo (Google DeepMind, 2024), and MovieDiffusion (Polyak et al., 2024), have achieved impressive visual realism and long-horizon coherence, motivating their use as foundations for generative world modeling. Recent world model systems further extend these capabilities toward persistent and interactive environments, either by explicitly constructing 3D scenes or by augmenting video generation with memory and feedback mechanisms (Chen et al., 2025d; Agarwal and others, 2025; Yu et al., 2024a; Wang and others, 2025; Team and others, 2025). While these approaches emphasize scalability, global coherence, and interactivity, the underlying physical state is often implicit, making fine-grained physical consistency—such as local geometry–motion coupling and precise object interactions—difficult to enforce. In contrast, our work explicitly targets fine-grained physics as a core objective, complementing large-scale world modeling with geometry- and motion-consistent 4D representations.
@@
Recent works further scale up 4D datasets by exploiting in-the-wild videos and automatic annotation pipelines. DynamicVerse introduces a large-scale multimodal 4D dataset derived from internet videos, providing metric-scale geometry, camera parameters, instance-level masks, motion cues, and hierarchical textual descriptions across over 100K dynamic scenes (Wen et al., 2025b). Complementary datasets such as OmniWorld and Ego-Exo4D extend this direction by covering diverse environments, viewpoints, and sensor modalities, facilitating evaluation of general-purpose 4D world understanding models (Zhou et al., 2025; Grauman et al., 2023). Collectively, these datasets form the foundation for benchmarking models that jointly reason about geometry, motion, and semantics in dynamic real-world scenes.
@@ G.3 Benchmarks for Physics-Aware Video Generation
To facilitate standardized evaluation, several benchmarks have been proposed to assess general video generation quality. VBench (Huang and others, 2024) and EvalCrafter (Liu and others, 2024b) provide comprehensive evaluation frameworks covering dimensions such as visual fidelity, diversity, temporal coherence, and scalability. T2V-CompBench (Sun and others, 2024) further focuses on compositionality, testing whether generated videos faithfully follow complex and multi-object textual prompts. While these benchmarks are effective for measuring appearance-level and semantic alignment, they do not explicitly evaluate physical plausibility.
Recent studies highlight the importance of physical realism in video generation and propose dedicated physics-oriented benchmarks. VideoPhy (Bansal et al., 2024a) evaluates physical commonsense reasoning in generated videos, while PhyGenBench (Meng and others, 2024) assesses adherence to basic physical laws such as object permanence and gravity. DEVIL (Liao and others, 2024) emphasizes dynamic realism, focusing on motion continuity and temporal consistency, and Physics-IQ (Motamed et al., 2025b) further probes understanding of fundamental physical principles including fluid dynamics and thermodynamics. Despite these advances, existing benchmarks are often limited in scenario diversity or focus on a narrow set of physical laws.
Motivated by these limitations, recent efforts introduce substantially expanded benchmarks that evaluate video generation models across a broader range of physical phenomena, including both realistic and deliberately unphysical dynamics (Meng and others, 2025). Empirical results show that state-of-the-art models perform significantly worse under identical settings, revealing the increased difficulty and diagnostic value of physics-aware benchmarks and underscoring the need for models that reason about physical consistency beyond visual realism.
@@ G.4 Physics-Aware Generation
This line of work first generates static images or 3D assets using generative models, and subsequently applies external physics simulators to animate them into dynamic scenes or videos. Representative approaches employ particle- or continuum-based simulators, such as the Material Point Method (MPM), to model object deformation and motion (Stomakhin et al., 2013a). Recent studies adopt this paradigm to produce physically plausible dynamics by post-processing generated content with physics engines (Lin and others, 2024a; Hsu and others, 2024). While this approach offers explicit and controllable physical behavior, it typically requires substantial manual effort and task-specific simulator design, limiting scalability and generalization.
@@
Another category of methods first performs physical simulation to obtain plausible motion trajectories or object states, which are then used as conditional inputs for video generation models. For example, PhysGen (Lin and others, 2024a) segments dynamic objects from an input image, simulates their motion using Newtonian mechanics, and conditions a video generator on both simulated trajectories and static backgrounds. Although this strategy improves controllability and motion realism, the generative model itself does not possess intrinsic physical reasoning capability, and physical rules and parameters must be manually specified for each scenario.
@@
More recent approaches aim to incorporate physics priors directly into the generative process by leveraging learned or pretrained models (Chang et al., 2024; Zhang and others, 2025; Chefer and others, 2025). These methods encode physical knowledge as latent dynamics, auxiliary supervision, or physics-aware evaluators. For instance, PhysT2V (Xue et al., 2025) employs large language models and vision–language models to iteratively evaluate and refine generated videos for improved physical consistency. However, such approaches often rely on implicit physical understanding acquired through data fitting, which may fail under physically challenging or out-of-distribution scenarios. Our method broadly follows this paradigm while grounding learned physics priors with physics-clean supervision to improve robustness and generalization.

### 06 Mar 2026 v2

@@ 1 Introduction
Recent video generation models (OpenAI, 2024; Google DeepMind, 2024; Polyak et al., 2024) and current generative world models(Chen et al., 2025c; Yu et al., 2024, 2023; Huang et al., 2025b) have achieved impressive visual realism and large-scale physics. Yet their success does not imply physical understanding: most are trained to match appearances rather than to model underlying physical state and dynamics. As a result, they frequently fail at fine-grained physics, producing locally inconsistent geometry, unstable object motion, and non-causal dynamics over time (Singer et al., 2022; Zhou et al., 2024). A faithful world model should instead represent how the environment evolves, requiring physics-consistent 4D representations.
A key challenge in physics-consistent 4D world modeling is the lack of scalable supervision for fine-grained physical dynamics. Learning such models requires dense, temporally aligned signals of geometry and motion, which are extremely difficult to obtain from real-world data at scale. Existing 4D datasets(Wen et al., 2025; Zhai et al., 2024) therefore tend to be appearance-rich but physics-poor, often dominated by camera motion. To address this limitation, we leverage the diversity and controllability of physics-based simulation to obtain large amounts of dynamic video data with accurate ground-truth geometry and motion, providing supervision that is otherwise inaccessible.
@@
To evaluate whether a model behaves as a physics-consistent 4D world model, we introduce a comprehensive evaluation protocol that combines content-level physics assessment and 4D world-level physics analysis. Specifically, in addition to established video-based physics benchmarks such as PhysicsIQ(Motamed et al., 2025), which evaluate physical plausibility from visual and semantic cues, we propose a 4D evaluation suite that probes geometry and motion world representations through trajectory-based analysis. Together, these complementary evaluations assess physical plausibility from what a model shows to how the underlying world evolves, providing a holistic diagnosis of physics-consistent 4D modeling beyond appearance and reconstruction quality.
@@ 2 Related Work
Recent video generation models(OpenAI, 2024; Google DeepMind, 2024) achieve strong visual realism and temporal coherence, but are largely optimized for appearance rather than physical world modeling. Several works attempt to introduce physical cues into video generation, such as modeling object dynamics(Zhang et al., 2024; Aira et al., 2024) or action-conditioned motion(Wang et al., 2025; Yuan et al., 2025), and a number of benchmarks(Motamed et al., 2025; Meng et al., 2025; Bansal et al., 2024) evaluate physical plausibility based on visual or semantic signals. While these approaches can improve or assess appearance-level physical realism, they operate primarily in image space and do not evaluate whether a model maintains coherent geometry, motion, or a consistent physical state over time. In contrast, our work focuses on physics consistency at the level of a 4D world, motivating geometry- and time-aware evaluation beyond appearance-centric video physics assessment. Detailed discussion please refer to Appendix G.1 G.2 G.3
@@
A growing body of work explores incorporating physical priors into generative models, including post hoc physical simulation after generation (Stomakhin et al., 2013; Chen et al., 2025a; Xie et al., 2024), simulation-guided generation with predefined dynamics (Yuan et al., 2023; Liu et al., 2024; Li et al., 2025), and learning physical priors from data (Li et al., 2024; Xu et al., 2024; Yang et al., 2025a). These methods typically focus on object-level or short-horizon dynamics, rely on specialized 3D representations, or remain optimization-based, limiting scalability and generalization. In contrast, our work leverages pretrained video diffusion models and treats fine-grained physics as a property of an evolving 4D world. For a detailed discussion, please refer to Appendix G.4
@@ 4 Simulation Data Collection
We present a physics-based simulation pipeline designed to address the limited physical fidelity and diversity of existing training data for 4D world modeling. In contrast to prior datasets that focus on narrow physical regimes or specific simulation paradigms (Huang et al., 2025a; Wang et al., 2025), Phys4D supports large-scale data generation with diverse object types, rich physical interactions, and accurate geometry and motion supervision. Figure 2 provides an overview of the simulation pipeline and the range of supported object categories.
@@ 6.1 Video-Based Physics Evaluation
We evaluate on Physics-IQ (Motamed et al., 2025), a benchmark designed to assess physical understanding in generative video models. The benchmark comprises 198 test scenarios covering diverse physical phenomena. Each scenario includes three camera viewpoints and two natural variations, yielding 1,188 test videos.
@@
We apply Phys4D to three backbone: WAN2.2-5b (Wan et al., 2025), CogVideoX-5b (Yang et al., 2025b), and Open-Sora V1.2 (Zheng et al., 2024b). For evaluation, we generate 5-second videos conditioned on the switch frame following the Physics-IQ protocol.
@@ F.1 Flow Matching for Video Generation
which induces a Gaussian conditional distribution $\mathbf{z}_{t}\sim\mathcal{N}(t\mathbf{z}_{1},(1-t)^{2}\mathbf{I})$. Notably, flow matching is not restricted to Gaussian endpoints or linear paths and can be generalized to interpolate between arbitrary distributions (Albergo and Vanden-Eijnden, 2023; Liu et al., 2023). This flexibility has recently enabled efficient large-scale generative modeling, including applications to high-dimensional and structured data such as videos, where flow-based formulations offer favorable trade-offs between sampling efficiency and temporal coherence (Xu et al., 2022; Lin and others, 2024; Zheng et al., 2024a).
@@ F.4 Policy Optimization with PPO
where $\hat{A}_{t}$ denotes the estimated advantage. This objective stabilizes policy updates and has been shown effective for fine-tuning diffusion- and flow-based generative models (Black and others, 2023; Ren and others, 2024; Liu et al., 2025).
@@ G.1 Large-Scale Video Generation and World Models
Recent large-scale video generation models, such as Sora (OpenAI, 2024), Veo (Google DeepMind, 2024), and MovieDiffusion (Polyak et al., 2024), have achieved impressive visual realism and long-horizon coherence, motivating their use as foundations for generative world modeling. Recent world model systems further extend these capabilities toward persistent and interactive environments, either by explicitly constructing 3D scenes or by augmenting video generation with memory and feedback mechanisms (Chen et al., 2025c; Agarwal and others, 2025; Yu et al., 2024; Huang et al., 2025b; Team and others, 2025). While these approaches emphasize scalability, global coherence, and interactivity, the underlying physical state is often implicit, making fine-grained physical consistency—such as local geometry–motion coupling and precise object interactions—difficult to enforce. In contrast, our work explicitly targets fine-grained physics as a core objective, complementing large-scale world modeling with geometry- and motion-consistent 4D representations.
@@
Recent works further scale up 4D datasets by exploiting in-the-wild videos and automatic annotation pipelines. DynamicVerse introduces a large-scale multimodal 4D dataset derived from internet videos, providing metric-scale geometry, camera parameters, instance-level masks, motion cues, and hierarchical textual descriptions across over 100K dynamic scenes (Wen et al., 2025). Complementary datasets such as OmniWorld and Ego-Exo4D extend this direction by covering diverse environments, viewpoints, and sensor modalities, facilitating evaluation of general-purpose 4D world understanding models (Zhou et al., 2025; Grauman et al., 2023). Collectively, these datasets form the foundation for benchmarking models that jointly reason about geometry, motion, and semantics in dynamic real-world scenes.
@@ G.3 Benchmarks for Physics-Aware Video Generation
To facilitate standardized evaluation, several benchmarks have been proposed to assess general video generation quality. VBench (Huang and others, 2024) and EvalCrafter (Liu and others, 2024) provide comprehensive evaluation frameworks covering dimensions such as visual fidelity, diversity, temporal coherence, and scalability. T2V-CompBench (Sun and others, 2024) further focuses on compositionality, testing whether generated videos faithfully follow complex and multi-object textual prompts. While these benchmarks are effective for measuring appearance-level and semantic alignment, they do not explicitly evaluate physical plausibility.
Recent studies highlight the importance of physical realism in video generation and propose dedicated physics-oriented benchmarks. VideoPhy (Bansal et al., 2024) evaluates physical commonsense reasoning in generated videos, while PhyGenBench (Meng and others, 2024) assesses adherence to basic physical laws such as object permanence and gravity. DEVIL (Liao and others, 2024) emphasizes dynamic realism, focusing on motion continuity and temporal consistency, and Physics-IQ (Motamed et al., 2025) further probes understanding of fundamental physical principles including fluid dynamics and thermodynamics. Despite these advances, existing benchmarks are often limited in scenario diversity or focus on a narrow set of physical laws.
Motivated by these limitations, recent efforts introduce substantially expanded benchmarks that evaluate video generation models across a broader range of physical phenomena, including both realistic and deliberately unphysical dynamics. Empirical results show that state-of-the-art models perform significantly worse under identical settings, revealing the increased difficulty and diagnostic value of physics-aware benchmarks and underscoring the need for models that reason about physical consistency beyond visual realism.
@@ G.4 Physics-Aware Generation
This line of work first generates static images or 3D assets using generative models, and subsequently applies external physics simulators to animate them into dynamic scenes or videos. Representative approaches employ particle- or continuum-based simulators, such as the Material Point Method (MPM), to model object deformation and motion (Stomakhin et al., 2013). Recent studies adopt this paradigm to produce physically plausible dynamics by post-processing generated content with physics engines. While this approach offers explicit and controllable physical behavior, it typically requires substantial manual effort and task-specific simulator design, limiting scalability and generalization.
@@
Another category of methods first performs physical simulation to obtain plausible motion trajectories or object states, which are then used as conditional inputs for video generation models. For example, PhysGen (Liu et al., 2024) segments dynamic objects from an input image, simulates their motion using Newtonian mechanics, and conditions a video generator on both simulated trajectories and static backgrounds. Although this strategy improves controllability and motion realism, the generative model itself does not possess intrinsic physical reasoning capability, and physical rules and parameters must be manually specified for each scenario.
@@
More recent approaches aim to incorporate physics priors directly into the generative process by leveraging learned or pretrained models (Wang et al., 2025; Zhang and others, 2025). These methods encode physical knowledge as latent dynamics, auxiliary supervision, or physics-aware evaluators. For instance, PhysT2V (Xue et al., 2025) employs large language models and vision–language models to iteratively evaluate and refine generated videos for improved physical consistency. However, such approaches often rely on implicit physical understanding acquired through data fitting, which may fail under physically challenging or out-of-distribution scenarios. Our method broadly follows this paradigm while grounding learned physics priors with physics-clean supervision to improve robustness and generalization.